The perimeter of a rectangle is 34 cm. The lenght of the rectangle is 7 cm more than the width. Find the dimensions of the rectangle

Answer by mananth(16949)

Let width be x
length = x+7
Area = 2(L+W)
2(x+x+7)=34
2(2x+7)=34
/2
2x+7=17
2x=17-7
2x=10
x=5cm width
length = x+7 = 12 cm
..
m.ananth@hotmail.ca